OpenStudy (anonymous):

helppppp..... how many times SMALLER is the 5 in tens place than the 5 in the thousands place?

OpenStudy (anonymous):

value of 5 in tens place is 5*10 = 50 value of 5 in thousands place is 5*1000 = 5000 Also 5000/100 = 50 So, it is 100 times smaller

OpenStudy (anonymous):

thank you so much!!! no one would help me :'( thanks!!!